OpenStudy (anonymous):

Write the rate as a unit rate. Peter Saltwood can paint 130 doors in 7 hours, while Carl Kepler can paint 53 doors in 4 hours. Who can paint doors faster, Peter or Carl? A. Peter B. Carl

OpenStudy (anonymous):

Carl

OpenStudy (anonymous):

P = 130/ 7 = 18 doors/ hr C = 53/ 4 = 13 doors/ hr => Peter is faster

OpenStudy (anonymous):

A. Peter